# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 9.3. Factor models
The proof of Theorem 4.6 for the factor model of Section 2.3 follows the same outline as in the previous two sections. However, an analogue of Propositions 9.1 and 9.6 becomes more complicated in this case.
@@
where $\mathbf{u}^{*}$ and $\mathbf{u}_{i}$ are $N$–dimensional unit vectors; $\mathbf{u}_{i}$, $1\leq i\leq N$, are pairwise orthogonal; $\mathbf{v}^{*}$ and $\mathbf{v}_{i}$ are $S$–dimensional unit vectors; $\mathbf{v}_{i}$, $1\leq i\leq N$, are pairwise orthogonal. If $a$ is a non-zero squared singular value of $X/\sqrt{S}$, i.e., an eigenvalue of $\frac{1}{S}XX^{\mathsf{T}}$, then either it solves an equation:
@@
[…]
@@
A somewhat similar statement can be found in Benaych-Georges and Nadakuditi (2012, Lemma 4.1). Also if for some $i$ we have $\lambda_{i}=0$, $\mathbf{u}^{*}=\mathbf{u}_{i}$, and $\langle\mathbf{u}^{*},\mathbf{u}_{i^{\prime}}\rangle=0$ for all $i\neq i^{\prime}$, then (9.29) turns into (9.18) of the previous subsection.
@@
We only describe the situation when all $\lambda_{i}$ are distinct and for each index $i$ either $\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle\neq 0$ or $\langle\mathbf{v}^{*},\mathbf{v}_{i}\rangle\neq 0$. Other cases can be obtained by continuously deforming the parameters.
@@
We should further identity the values of $\tilde{\alpha}$ and $\tilde{\beta}$, for which we plug (9.30) back into their definitions, getting a system of two equations:
@@
An analogue of Corollaries 9.3 and 9.8 becomes more delicate for the factor model, which, however, does not lead to any significant changes in the ways we use it in our arguments.
@@
The determinant of this matrix is $(c^{2}-d)\theta S<0$, because $\mathbf{u}^{*}$ is a unit vector. Hence, the matrix has one positive and one negative eigenvalues. It remains to use Corollary 9.3 twice. ∎
Next we state an analogue of Lemma 9.4.
@@
One can replace the vectors $\mathbf{u}_{i}^{*}$ and $\mathbf{v}_{i}^{*}$ in the factor model ${X=\sum_{i=1}^{r}\sqrt{\theta_{i}}\sqrt{S}\cdot\mathbf{u}^{*}_{i}(\mathbf{v}^{*}_{i})^{\mathsf{T}}+\mathcal{E}}$ of (2.10) with $2r$ vectors uniformly distributed on the $N$–dimensional and $S$–dimensional unit spheres, respectively, independent from each other and from $\mathcal{E}$: Theorem 4.6 for deterministic orthonormal vectors as in Section 2.3 is equivalent to the same theorem for independent unit vectors.
@@
We first note that the asymptotics in Theorem 4.6 does not depend on the choice of the unit vectors $\{\mathbf{u}_{i}^{*}\}_{i=1}^{r}$ and $\{\mathbf{v}_{i}^{*}\}_{i=1}^{r}$ in (2.10) as long as they form two orthonormal systems. Indeed, any $r$ orthonormal vectors can be obtained from any other $r$ orthonormal vectors by an orthogonal transformation of the space. Such transformation does not change the eigenvalues of $\frac{1}{S}XX^{\mathsf{T}}$, and also does not change the probability distribution of the matrix $\mathcal{E}$ (which uses Gaussianity of its matrix elements).
Now let $\tilde{\mathbf{u}}_{1}^{*},\dots,\tilde{\mathbf{u}}_{r}^{*}$ be $r$ independent vectors uniformly distributed on the $N$–dimensional unit sphere and let $\tilde{\mathbf{v}}_{1}^{*},\dots,\tilde{\mathbf{v}}_{r}^{*}$ be $r$ independent vectors uniformly distributed on the $S$–dimensional unit sphere. We consider the matrix $M_{r}=\sum_{i=1}^{r}\sqrt{\theta_{i}}\sqrt{S}\tilde{\mathbf{u}}_{i}^{*}(\tilde{\mathbf{v}}_{i}^{*})^{\mathsf{T}}$ and would like to decompose it as $M_{r}=\sum_{i=1}^{r}\sqrt{\theta_{i}^{\prime}}{\sqrt{S}}\mathbf{u}_{i}^{*}(\mathbf{v}_{i}^{*})^{\mathsf{T}}$ with orthonormal vectors $\mathbf{u}_{i}^{*}$ and $\mathbf{v}_{i}^{*}$; $\tilde{\mathbf{u}}_{i}^{*}$ and $\tilde{\mathbf{v}}_{i}^{*}$ were not orthonormal. Clearly, $\sqrt{\theta_{i}^{\prime}}$ are non-zero singular values of $M_{r}$ and $\mathbf{u}_{i}^{*}$, $\mathbf{v}_{i}^{*}$ are corresponding left and right singular vectors. We claim that
@@
where $(\mathbf{u}_{i},\mathbf{v}_{i},\theta^{\prime}_{i})$ are singular vectors and values of $M_{r-1}$ for $1\leq i\leq{r-1}$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ with $i\geq r$ complement those to othonormal bases of $N$– and $S$–dimensional spaces, respectively. We note that after multiplying by the denominators $a$, $a-\theta^{\prime}_{1}$, …, $a-\theta^{\prime}_{r-1}$, (9.33) becomes a polynomial equation of degree $r$ (there is a cancelation between the left and right hand sides which guarantees that denominators $(a-\theta^{\prime}_{i})^{2}$ do not appear) and therefore it has $r$ roots. Let us locate these roots.
@@
(9.35) $\lim_{N\to\infty}\sqrt{N}(\lambda_{i}-\lambda(\theta_{i}))$ =$\stackrel{{\scriptstyle d}}{{=}}\mathcal{N}(0,V(\theta_{i})),\qquad 1\leq i\leq q-1,$
(9.36) $\lim_{N\to\infty}N^{2/3}\frac{\lambda_{i}-(1+\gamma)^{2}}{\gamma(1+\gamma)^{4/3}}$ =$\stackrel{{\scriptstyle d}}{{=}}\mathfrak{a}_{i-q+1},\qquad i\geq q,$
where $\mathcal{N}(0,V(\theta_{i}))$, $10\leq i\leq q-$ 1, are independent and $\{\mathfrak{a}_{j}\}_{j\geq 1}$ are points of the Airy1 point process independent from (9.35). In addition, (8.46) holds for the eigenvalues of $\frac{1}{S}XX^{\mathsf{T}}$ both with $h(x)=1$ and with $h(x)=\sqrt{x}$.
@@
[…]
which are the same as in (9.22), and the constants of (8.45) are computed to be $\mathfrak{s}=\frac{\sqrt{\lambda_{+}-\lambda_{-}}}{2\gamma^{2}\lambda_{+}}=\frac{1}{\gamma^{3/2}(1+\gamma)^{2}}$, $\mathfrak{m}=\frac{(1+\gamma)^{2}+\gamma^{2}-1}{2\gamma^{2}(1+\gamma)^{2}}=\frac{1}{\gamma(1+\gamma)}$. For the choice $h(x)=\sqrt{x}$, the corresponding function $m(z)$ can be also computed, but we do not need this expression.
Statements similar to (9.35) are known from Onatski (2012) and Benaych-Georges and Nadakuditi (2012); (9.36) is harder to locate in the literature (although it is also probably known to the specialists). Alternatively, the proof can be obtained by the same argument as in Proposition 9.5, by induction on $r$ with the base case $r=0$ given in Johnstone (2001); Soshnikov (2002) and the step based on Proposition 9.10. We only highlight the key computation, which is an analogue of (9.11) and (9.12).
@@
Using the Marchenko-Pastur law (see Theorem 8.24 or Bai and Silverstein (2010, Chapter 3 and Theorem 9.10)), formulas for the limit (9.22), and $\frac{N}{S}=\gamma^{2}+O\left(\frac{1}{N}\right)$ the equation (9.39) becomes:
(9.40) $1-a\theta\,m(a)\tilde{m}(a)+\sum_{i=1}^{N}\frac{\xi_{i}^{2}-1}{N}\left[1-\frac{a\theta}{a-\lambda_{i}}\frac{1}{S}\sum\limits_{j=1}^{S}\frac{1}{a-\lambda_{j}}\right]\\ +\sum_{j=1}^{S}\frac{\eta_{j}^{2}-1}{S}\left[1-\frac{a\theta}{a-\lambda_{j}}\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{a-\lambda_{i}}\right]-2\sqrt{\frac{\theta}{NS}}\sum\limits_{i=1}^{N}\frac{\sqrt{\lambda_{i}}\xi_{i}\eta_{i}}{a-\lambda_{i}}=o\left(\frac{1}{\sqrt{N}}\right),$
where we used the notation
$\tilde{m}(a)=\left[\gamma^{2}m(a)+\frac{1-\gamma^{2}}{a}\right].$
Since $\xi^{2}_{i}-1$, $\eta_{j}^{2}-1$, $\xi_{i}\eta_{i}$ are three mean 00 i.i.d. in $i$ sequences, we are in a position to apply the Central Limit Theorem; note that all pairwise covariances between these random variables vanish (using $\mathbb{E}(\xi_{i}^{2}-1)\xi_{i}\eta_{i}=0$), and therefore different sums give rise to independent Gaussian limits. Using $\mathbb{E}(\xi_{i}^{2}-1)^{2}=\mathbb{E}(\xi_{j}^{2}-1)^{2}=2$, $\mathbb{E}\xi_{i}^{2}\eta_{i}^{2}=1$ and applying CLT, we further transform (9.40) into
@@
In the leading order, the solution $a$ is found from the equation $01-a\theta\,m(a)\tilde{m}(a)=$ 0. Plugging the formula for $m(z)$ from (9.22), this becomes
@@
In addition, the expression after $\mathcal{N}(0,1)$ under $[\cdot]^{1/2}$ in (9.42) simplifies upon plugging $a$ from (9.44): using (9.45), (9.46), we get
@@
Its solution gives the desired asymptotic statement (9.35), matching the formula for $V(\theta)$ in (2.12). ∎
Proof of Theorem 4.6 for the factor model of Section 2.3.
The constants (4.8) simplify to $\kappa_{1}=\gamma(1+\gamma)^{4/3}$, $\kappa_{2}=\frac{1}{\gamma(1+\gamma)^{2/3}}$ – exactly the same as in the proofs for the spiked covariance model in Section 2.2. Hence, Assumption 8.19 will be satisfied.
@@
where the asymptotics of $(\lambda_{i})_{i=1}^{N}$ is given to us by Proposition 9.14 and $\xi_{i}$ and $\eta_{j}$ are i.i.d. $\mathcal{N}(0,1)$. Arguing as in the previous sections, the $q-1$ largest roots of the equation are close to $\lambda_{1},\dots,\lambda_{q-1}$, resulting in (4.6). In order to establish (4.7), we need to approximate (9.48) for $a$ close to $\lambda_{+}=(1+\gamma)^{2}$ and locate the root of the equation in the $(\lambda_{q},\lambda_{q-1})$ interval. We change the variables
@@
recall that $\theta_{c}=\gamma$, $\theta_{q}=\gamma+N^{-1/3}\tilde{\theta}$, and aim to apply Theorem 8.20 to (9.48). For this computation, we can approximate $\sum_{i=1}^{N}\xi_{i}^{2}\approx N$ and $\sum_{j=1}^{S}\eta_{j}^{2}\approx S$, because the relative errors in these approximations are of order $N^{-1/2}$, which is smaller than $N^{-1/3}$, the scale of our interest. Hence, (9.48) becomes
@@
For the factors in the right-hand side of (9.50) we apply Theorem 8.20 with $h(x)=1$. For the left-hand side we write
@@
(9.52) $-\tilde{\theta}\frac{1}{\gamma}+o(1)=\frac{1}{(1+\gamma)^{1/3}}\mathcal{G}^{(3)}(b)+\frac{\gamma}{(1+\gamma)^{1/3}}\mathcal{G}^{(4)}(b)+\gamma^{1/2}\frac{\left(\mathcal{G}^{(1)}(b)-\mathcal{G}^{(2)}(b)\right)}{(1+\gamma)^{1/3}}.$
At this step an algebraic miracle happens; the following claim is responsible for the appearance of exactly the same function $\mathcal{T}(\Theta)$ in the asymptotics of the factor model:
@@
The claim is established by recalling which noises enter into the functions $\mathcal{G}^{(k)}$, $k=1,2,3,4$: these are $\left(\frac{\xi_{j}+\eta_{j}}{\sqrt{2}}\right)^{2}$, $\left(\frac{\xi_{j}-\eta_{j}}{\sqrt{2}}\right)^{2}$, $\xi_{j}^{2}$, $\eta_{j}^{2}$, respectively. Hence, in the linear combination of $\mathcal{G}^{(k)}$ of (9.52), the noises combine into
@@
Using the claim, we conclude that (in distribution) the solution of the equation (9.52) converges as $N\to\infty$ to the solution of
(9.53) $-\tilde{\theta}\frac{1}{\gamma(1+\gamma)^{2/3}}=\mathcal{G}(b).$
Recognizing the constant $\kappa_{2}$ in the left-hand side of the last formula and comparing with Definition 4.2, we arrive at (4.7). ∎

### v2

@@ 9.3. Factor models
An analogue of Propositions 9.1 and 9.6 is more complicated here.
@@
where $\mathbf{u}^{*}$ and $\mathbf{u}_{i}$ are NN–dimensional unit vectors; $\mathbf{u}_{i}$, $1\leq i\leq N$, are pairwise orthogonal; $\mathbf{v}^{*}$ and $\mathbf{v}_{i}$ are SS–dimensional unit vectors; $\mathbf{v}_{i}$, $1\leq i\leq N$, are pairwise orthogonal. If $a$ is a non-zero squared singular value of $X/\sqrt{S}$, i.e., an eigenvalue of $\frac{1}{S}XX^{\mathsf{T}}$, then either it solves
@@
where $\mathbf{v}_{N+1}$, $\mathbf{v}_{S}$ are arbitrary vectors complimenting $\mathbf{v}_{1},\dots,\mathbf{v}_{N}$ to an orthonormal basis, $\langle\cdot,\cdot\rangle$ is the scalar product, and we set $\lambda_{j}=0$ for $N<j\leq S$. Or $a=\lambda_{i}$ for $1\leq i\leq N$, where $\lambda_{i}$ has multiplicity one, $\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle=\langle\mathbf{v}^{*},\mathbf{v}_{i}\rangle=0$, and the equation (9.30) holds with the $i$-th terms excluded; or $a=\lambda_{i}$, where $\lambda_{i}$ has multiplicity larger than one.
@@
A related statement is Benaych-Georges and Nadakuditi (2012, Lemma 4.1). If for some $i$, $\lambda_{i}=0$, $\mathbf{u}^{*}=\mathbf{u}_{i}$, $\langle\mathbf{u}^{*},\mathbf{u}_{i^{\prime}}\rangle=0$ for all $i\neq i^{\prime}$, then (9.30) turns into (9.18).
@@
We only deal with distinct $\lambda_{i}$ and assuming for each index $i$ either $\langle\mathbf{u}^{*},\mathbf{u}_{i}\rangle\neq 0$ or $\langle\mathbf{v}^{*},\mathbf{v}_{i}\rangle\neq 0$. Other cases are obtained by continuous deformations.
@@
We find the values of $\tilde{\alpha}$ and $\tilde{\beta}$, by plugging (9.31) back into their definitions, getting:
@@
This matrix has one positive and one negative eigenvalue. because its determinant is $(c^{2}-d)\theta S<0$, since $\mathbf{u}^{*}$ is a unit vector. It remains to use Corollary 9.3 twice. ∎
@@
In the factor model $X=\sum_{i=1}^{r}\sqrt{\theta_{i}}\sqrt{S}\cdot\mathbf{u}^{*}_{i}(\mathbf{v}^{*}_{i})^{\mathsf{T}}+\mathcal{E}$ of (2.10), replacing $\mathbf{u}_{i}^{*}$ and $\mathbf{v}_{i}^{*}$ with $2r$ independent vectors uniformly distributed on the NN– and SS–dimensional unit spheres, respectively (independent of each other and $\mathcal{E}$) leaves Theorem 4.6 unchanged.
@@
The asymptotics in Theorem 4.6 does not depend on the choice of $\{\mathbf{u}_{i}^{*}\}_{i=1}^{r}$ and $\{\mathbf{v}_{i}^{*}\}_{i=1}^{r}$ in (2.10) as long as they form two orthonormal systems. Indeed, any $r$ orthonormal vectors can be obtained from any other $r$ orthonormal vectors by an orthogonal transformation. Such transformation does not change the eigenvalues of $\frac{1}{S}XX^{\mathsf{T}}$, and also does not change the probability distribution of the matrix $\mathcal{E}$ (which uses Gaussianity of its matrix elements).
Now let $\tilde{\mathbf{u}}_{1}^{*},\dots,\tilde{\mathbf{u}}_{r}^{*}$ be $r$ independent vectors uniformly distributed on the NN–dimensional unit sphere and let $\tilde{\mathbf{v}}_{1}^{*},\dots,\tilde{\mathbf{v}}_{r}^{*}$ be $r$ independent vectors uniformly distributed on the SS–dimensional unit sphere. We consider the matrix $M_{r}=\sum_{i=1}^{r}\sqrt{\theta_{i}}\sqrt{S}\tilde{\mathbf{u}}_{i}^{*}(\tilde{\mathbf{v}}_{i}^{*})^{\mathsf{T}}$ and would like to decompose it as $M_{r}=\sum_{i=1}^{r}\sqrt{\theta_{i}^{\prime}}{\sqrt{S}}\mathbf{u}_{i}^{*}(\mathbf{v}_{i}^{*})^{\mathsf{T}}$ with orthonormal vectors $\mathbf{u}_{i}^{*}$ and $\mathbf{v}_{i}^{*}$; $\tilde{\mathbf{u}}_{i}^{*}$ and $\tilde{\mathbf{v}}_{i}^{*}$ were not orthonormal. Clearly, $\sqrt{\theta_{i}^{\prime}}$ are non-zero singular values of $M_{r}$ and $\mathbf{u}_{i}^{*}$, $\mathbf{v}_{i}^{*}$ are corresponding left and right singular vectors. We claim that
@@
where $(\mathbf{u}_{i},\mathbf{v}_{i},\theta^{\prime}_{i})$ are singular vectors and values of $M_{r-1}$ for $1\leq i\leq{r-1}$, and $\mathbf{u}_{i}$, $\mathbf{v}_{i}$ with $i\geq r$ complement those to othonormal bases of NN– and SS–dimensional spaces, respectively. We note that after multiplying by the denominators $a$, $a-\theta^{\prime}_{1}$, …, $a-\theta^{\prime}_{r-1}$, (9.34) becomes a polynomial equation of degree $r$ (there is a cancelation between the left and right hand sides which guarantees that denominators $(a-\theta^{\prime}_{i})^{2}$ do not appear) and therefore it has $r$ roots. Let us locate these roots.
@@
(9.37) $\lim_{N\to\infty}N^{2/3}\frac{\lambda_{i}-(1+\gamma)^{2}}{\gamma(1+\gamma)^{4/3}}$ $\stackrel{{\scriptstyle d}}{{=}}\mathfrak{a}_{i-q+1},\qquad i\geq q,$
where $\mathcal{N}(0,V(\theta_{i}))$ are independent over $i$ and with points of the Airy1 point process $\{\mathfrak{a}_{j}\}_{j\geq 1}$. In addition, (8.45) holds for the eigenvalues of $\frac{1}{S}XX^{\mathsf{T}}$ with $h(x)=1$ and with $h(x)=\sqrt{x}$.
@@
(9.38) $\mu(x)\mathrm{d}x=\frac{1}{2\pi}\frac{\sqrt{(\lambda_{+}-x)(x-\lambda_{-})}}{\gamma^{2}x}\,\mathbf{1}_{[\lambda_{-},\lambda_{+}]}\,\mathrm{d}x,\qquad m(z)=\frac{z+\gamma^{2}-1-\sqrt{(z-\lambda_{+})(z-\lambda_{-})}}{2\gamma^{2}z},$
which are the same as in (9.22), and the constants of (8.44) are computed to be $\mathfrak{s}=\frac{\sqrt{\lambda_{+}-\lambda_{-}}}{2\gamma^{2}\lambda_{+}}=\frac{1}{\gamma^{3/2}(1+\gamma)^{2}}$, $\mathfrak{m}=\frac{(1+\gamma)^{2}+\gamma^{2}-1}{2\gamma^{2}(1+\gamma)^{2}}=\frac{1}{\gamma(1+\gamma)}$.
Statements similar to (9.36) are known from Onatski (2012) and Benaych-Georges and Nadakuditi (2012); (9.37) is harder to locate in the literature. The proof can be obtained by the same argument as in Proposition 9.5, by induction on $r$ with the base case $r=0$ given in Johnstone (2001); Soshnikov (2002) and the step based on Proposition 9.11. We only highlight the key computation, which is an analogue of (9.11) and (9.12).
@@
Using the Marchenko-Pastur law (Theorem 8.24 or Bai and Silverstein (2010, Chapter 3 and Theorem 9.10)), formulas (9.22), the notation $\tilde{m}(a)=\left[\gamma^{2}m(a)+\frac{1-\gamma^{2}}{a}\right]$, and $\frac{N}{S}=\gamma^{2}+O\left(\frac{1}{N}\right)$ the equation (9.40) becomes:
(9.41) $1-a\theta\,m(a)\tilde{m}(a)+\sum_{i=1}^{N}\frac{\xi_{i}^{2}-1}{N}\left[1-\frac{a\theta}{a-\lambda_{i}}\frac{1}{S}\sum\limits_{j=1}^{S}\frac{1}{a-\lambda_{j}}\right]\\ +\sum_{j=1}^{S}\frac{\eta_{j}^{2}-1}{S}\left[1-\frac{a\theta}{a-\lambda_{j}}\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{a-\lambda_{i}}\right]-2\sqrt{\frac{\theta}{NS}}\sum\limits_{i=1}^{N}\frac{\sqrt{\lambda_{i}}\xi_{i}\eta_{i}}{a-\lambda_{i}}=o\left(\frac{1}{\sqrt{N}}\right).$
Since $\xi^{2}_{i}-1$, $\eta_{j}^{2}-1$, $\xi_{i}\eta_{i}$ are three mean 0 i.i.d. in $i$ sequences, we are in a position to apply the Central Limit Theorem; note that all pairwise covariances between these random variables vanish (using $\mathbb{E}(\xi_{i}^{2}-1)\xi_{i}\eta_{i}=0$), and therefore different sums give rise to independent Gaussian limits. Using $\mathbb{E}(\xi_{i}^{2}-1)^{2}=\mathbb{E}(\xi_{j}^{2}-1)^{2}=2$, $\mathbb{E}\xi_{i}^{2}\eta_{i}^{2}=1$ and applying CLT, we further transform (9.41) into
@@
In the leading order, the solution $a$ is found from the equation $1-a\theta\,m(a)\tilde{m}(a)=0$. Plugging the formula for $m(z)$ from (9.22), this becomes
@@
In addition, the expression after $\mathcal{N}(0,1)$ under $\cdot]^{1/2}$ in (9.43) simplifies upon plugging $a$ from (9.45): using (9.46), (9.47), we get
@@
Its solution gives the desired (9.36), matching the formula for $V(\theta)$ in (2.12). ∎
Proof of Theorem 4.6 for the factor model.
The constants (4.8) are $\kappa_{1}=\gamma(1+\gamma)^{4/3}$, $\kappa_{2}=\frac{1}{\gamma(1+\gamma)^{2/3}}$ – the same as for the spiked covariance model. Assumption 8.19 is satisfied.
@@
where the asymptotics of $(\lambda_{i})_{i=1}^{N}$ is given by Proposition 9.15 and $\xi_{i}$ and $\eta_{j}$ are i.i.d. $\mathcal{N}(0,1)$. As in the previous sections, the q1q-1 largest roots of the equation are close to $\lambda_{1},\dots,\lambda_{q-1}$, resulting in (4.6). To establish (4.7), we approximate (9.49) for $a$ close to $\lambda_{+}=(1+\gamma)^{2}$ and locate the root of the equation in $(\lambda_{q},\lambda_{q-1})$ interval. We change the variables
@@
recall that $\theta_{c}=\gamma$, $\theta_{q}=\gamma+N^{-1/3}\tilde{\theta}$, and aim to apply Theorem 8.20 to (9.49). We can approximate $\sum_{i=1}^{N}\xi_{i}^{2}\approx N$ and $\sum_{j=1}^{S}\eta_{j}^{2}\approx S$, because the relative errors of order $N^{-1/2}$ in these approximations are smaller than $N^{-1/3}$, the scale of our interest. (9.49) becomes
@@
In the right-hand side we apply Theorem 8.20 with $h(x)=1$. For the left-hand side we write
@@
(9.53) $\tilde{\theta}\frac{1}{\gamma}+o(1)=\frac{1}{(1+\gamma)^{1/3}}\mathcal{G}^{(3)}(b)+\frac{\gamma}{(1+\gamma)^{1/3}}\mathcal{G}^{(4)}(b)+\gamma^{1/2}\frac{\left(\mathcal{G}^{(1)}(b)-\mathcal{G}^{(2)}(b)\right)}{(1+\gamma)^{1/3}}.$
At this step an algebraic miracle happens, leading to the appearance of the familiar $\mathcal{T}(\Theta)$:
@@
For the proof, recall that noises entering into $\mathcal{G}^{(k)}$, $k=1,2,3,4$, are $\left(\frac{\xi_{j}+\eta_{j}}{\sqrt{2}}\right)^{2}$, $\left(\frac{\xi_{j}-\eta_{j}}{\sqrt{2}}\right)^{2}$, $\xi_{j}^{2}$, $\eta_{j}^{2}$, respectively. Hence, in the linear combination of $\mathcal{G}^{(k)}$ of (9.53), the noises combine into
@@
Using the claim, the solution of the equation (9.53) converges to the solution of $\tilde{\theta}\frac{1}{\gamma(1+\gamma)^{2/3}}=\mathcal{G}(b).$ Recognizing $\kappa_{2}$ and comparing with Definition 4.2, we arrive at (4.7). ∎